# Curiosity Driven Exploration to Optimize Structure-Property Learning in Microscopy

### v1

@@ III.3 Real-time microscope deployment
Given these promising results on pre-acquired data, we moved to implement the curiosity algorithm on the microscope for real-time adaptive sampling. For additional difficulty, we changed the sample to one with a more complex domain structure, a thin film $PbTiO_{3}$ sample with a hierarchical domain strcuture that has been previously investigated liu2022exploring , and implemented the method using our AEcroscopy platform for microscope automation and acquisitionliu2024aecroscopy . Here, we tested both the spec2Im as well as the inverse, Im2spec, for the curiosity algorithm, and plot the results in Fig. 7. The exploration path the algorithm took for the Spec2Im case is shown in Fig. 8(a), and indicates a diverse spread of points across multiple different domain structures. Predicted errors are still spatially localized, but observing examples of predicted images compared with ground truth images show a decent predictive capability (Fig. 8(c,d)). The Im2Spec model shows a different exploration path, with many more points in the darker regions of the image, and the error map appears highly localized, potentially indicating that more points would need to be measured for more accurate modeling. Nevertheless, analysis of the actual predictions shows a decent corroboration with the ground truth (e.g., Fig. 8(g)).
@@ V Data Availability Statement
The code and data used for model training and analysis can be found in the repository: https://github.com/cylindrical-penguin/Curiosity-Driven-RL-for-PFM.
@@ VII Acknowledgment
Algorithmic development was supported by the US Department of Energy, Office of Science, Office of Basic Energy Sciences, MLExchange Project, award number 107514. The experimental work was supported by the Center for Nanophase Materials Sciences (CNMS), which is a US Department of Energy, Office of Science User Facility at Oak Ridge National Laboratory. GN acknowledges support from QIS Infrastructure Project. HF acknowledges MEXT Program: Data Creation and Utilization Type Material Research and Development Project (Grant Number JPMXP1122683430). J.-C.Y. acknowledges the financial support from the National Science and Technology Council (NSTC), Taiwan, under grant numbers NSTC 112-2112-M-006-020-MY3 and 113-2124-M-006-010.

### v2

@@ III.3 Real-time microscope deployment
Given these promising results on pre-acquired data, we moved to implement the curiosity algorithm on the microscope for real-time adaptive sampling. For additional difficulty, we changed the sample to one with a more complex domain structure, a thin film $PbTiO_{3}$ sample with a hierarchical domain structure that has been previously investigated liu2022exploring , and implemented the method using our AEcroscopy platform for microscope automation and acquisitionliu2024aecroscopy . Here, we tested both the spec2Im as well as the inverse, Im2spec, for the curiosity algorithm, and plot the results in Fig. 8. The exploration path the algorithm took for the Spec2Im case is shown in Fig. 8(a), and indicates a diverse spread of points across multiple different domain structures. Predicted errors are still spatially localized, but observing examples of predicted images compared with ground truth images show a decent predictive capability (Fig. 8(c,d)). The Im2Spec model shows a different exploration path, with many more points in the darker regions of the image, and the error map appears highly localized, potentially indicating that more points would need to be measured for more accurate modeling. Nevertheless, analysis of the actual predictions shows a decent corroboration with the ground truth (e.g., Fig. 8(g)).
@@ V Data Availability Statement
The data and the code used in this work are available at Zenedo (doi: 10.5281/zenodo.15777800). The Jupyter notebooks can also be found on: https://github.com/cylindrical-penguin/Curiosity-Driven-RL-for-PFM. The scripts presented here utilize open-source repositories im2spec(github.com/ziatdinovmax/im2spec) and atomai (github.com/pycroscopy/atomai). The deployment of the workflows on the microscopes were interfaced using aecroscopy (https://github.com/yongtaoliu/aecroscopy.pyae).
@@ VII Acknowledgment
Algorithmic development was supported by the US Department of Energy, Office of Science, Office of Basic Energy Sciences, MLExchange Project, award number 107514. The experimental work was supported by the Center for Nanophase Materials Sciences (CNMS), which is a US Department of Energy, Office of Science User Facility at Oak Ridge National Laboratory. GN acknowledges support from QIS Infrastructure Project (FWP ERKCZ62): Precision Atomic Assembly for Quantum Information Science. H.F. was supported by MEXT Program: Data Creation and Utilization Type Material Research and Development Project (No. JPMXP1122683430) and MEXT Initiative to Establish Next-generation Novel Integrated Circuits Centers (X-NICS) (JPJ011438), and the Japan Science and Technology Agency (JST) as part of Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE), Grant Number JPMJAP2312. J.-C.Y. acknowledges the financial support from the National Science and Technology Council (NSTC), Taiwan, under grant numbers NSTC 112-2112-M-006-020-MY3 and 113-2124-M-006-010.